# Can Static Black Holes in Massive Gravity Serve as Candidates for Aschenbach-Like Phenomena?

### v1

@@ III.3 ModMax-dRGT-like Black hole in massive gravity
As can be seen in Fig. (10), since in this model the $\gamma$ parameter will have a significant effect on the gravitational behavior of the model, in such a way that small values (e.g. $\gamma<0.6686$) will push the model towards extremalization and larger values will lead to black hole behavior, we therefore plot the metric function in two cases.
@@
Again, with considering the Eq. (2) and Eq. (5) we have:
$H=\frac{\sqrt{9-\frac{9M}{r}-3r^{2}\Lambda+\frac{9q^{2}{\mathrm{e}}^{-\gamma}}{r^{2}}+9\left(\frac{rc_{1}}{2}+Cc_{2}\right)Cm_{g}^{2}}}{3\sin\!\left(\theta\right)r},$ (27)
$\phi^{r_{h}}=-\frac{\csc\!\left(\theta\right)\left(4C^{2}r^{2}c_{2}m_{g}^{2}+C\,r^{3}c_{1}m_{g}^{2}+8q^{2}{\mathrm{e}}^{-\gamma}+4r^{2}-6Mr\right)}{4r^{4}},$ (28)
$\phi^{\Theta}=-\frac{\sqrt{6}\,\sqrt{\frac{3C\,r^{3}c_{1}m_{g}^{2}-2r^{4}\Lambda+\left(6C^{2}c_{2}m_{g}^{2}+6\right)r^{2}-6Mr+6q^{2}{\mathrm{e}}^{-\gamma}}{r^{2}}}\,\cot\!\left(\theta\right)\csc\!\left(\theta\right)}{6r^{2}}.$ (29)
Figure 13: Fig (13a): The photon spheres location at $(r,\theta)=(2.0913,1.57)$ and $(r,\theta)=(4.4789,1.57)$ with respect to $M=2,\Lambda=-0.5,\gamma=2,q=1,C=0.4,c_{1}=-12,c_{2}=25,m_{g}=0.5$ in the $(r-\theta)$ plane of the normal vector field $n$ , (13b): the topological potential H(r) for the ModMax-dRGT-like Black hole in massive gravity
@@
Figure 14: Fig (14a): $\beta$ diagram (14b):MSCO localization and space classification for the ModMax-dRGT-like Black hole in massive gravity
@@ IV Conclusions
This motivated investigations into non-rotating black hole models to explore whether such behavior could emerge independent of frame dragging 14 ; 18 ; 19 .
In our previous study 14 , we examined the mechanism responsible for this phenomenon across various ultra-compact gravitational configurations, both in the form of black holes and naked singularities. We found a consistent feature shared by all examined models 14 ; 18 ; 19 : the presence of a stable gravitational potential minimum—either local or global—which corresponds to a stable photon sphere. In naked singularity models, the absence of an event horizon naturally permits such minima. However, in black hole configurations, the existence of a stable minimum depends on its placement outside the event horizon, adding a geometric constraint.
@@
Following the analysis of the 4D Einstein–Gauss–Bonnet Massive Gravity model, alongside results reported in 19 , we posed a new question:

### v3

@@ 3.3 ModMax-dRGT-like Black hole in massive gravity
Again, with considering the Eq. (2.2) and Eq. (2.5) we have:
$H=\frac{\sqrt{9-\frac{9M}{r}-3r^{2}\Lambda+\frac{9q^{2}{\mathrm{e}}^{-\gamma}}{r^{2}}+9\left(\frac{rc_{1}}{2}+Cc_{2}\right)Cm_{g}^{2}}}{3\sin\!\left(\theta\right)r},$ (3.16)
$\phi^{r_{h}}=-\frac{\csc\!\left(\theta\right)\left(4C^{2}r^{2}c_{2}m_{g}^{2}+C\,r^{3}c_{1}m_{g}^{2}+8q^{2}{\mathrm{e}}^{-\gamma}+4r^{2}-6Mr\right)}{4r^{4}},$ (3.17)
$\phi^{\Theta}=-\frac{\sqrt{6}\,\sqrt{\frac{3C\,r^{3}c_{1}m_{g}^{2}-2r^{4}\Lambda+\left(6C^{2}c_{2}m_{g}^{2}+6\right)r^{2}-6Mr+6q^{2}{\mathrm{e}}^{-\gamma}}{r^{2}}}\,\cot\!\left(\theta\right)\csc\!\left(\theta\right)}{6r^{2}}.$ (3.18)
Figure 13: (13a): The photon spheres location at $(r,\theta)=(2.0913,1.57)$ and $(r,\theta)=(4.4789,1.57)$ with respect to $M=2,\Lambda=-0.5,\gamma=2,q=1,C=0.4,c_{1}=-12,c_{2}=25,m_{g}=0.5$ in the $(r-\theta)$ plane of the normal vector field $n$ , (13b): the topological potential H(r) for the ModMax-dRGT-like Black hole in massive gravity
@@
Figure 14: (14a): $\beta$ diagram (14b):MSCO localization and space classification for the ModMax-dRGT-like Black hole in massive gravity
@@ 4 Conclusions
This motivated investigations into non-rotating black hole models to explore whether such behavior could emerge independent of frame dragging [11, 15, 16].
In our previous study [11], we examined the mechanism responsible for this phenomenon across various ultra-compact gravitational configurations, both in the form of black holes and naked singularities. We found a consistent feature shared by all examined models [11, 15, 16]: the presence of a stable gravitational potential minimum—either local or global—which corresponds to a stable photon sphere. In naked singularity models, the absence of an event horizon naturally permits such minima. However, in black hole configurations, the existence of a stable minimum depends on its placement outside the event horizon, adding a geometric constraint.
@@
Following the analysis of the 4D Einstein–Gauss–Bonnet Massive Gravity model, alongside results reported in [16], we posed a new question: